# On a Roll Again: Analysis of a Dice Removal Game

### v2

@@ Abstract.
We thank all of Polymath Jr’s staff, whose work has been key to the writing of this paper. In particular, we extend our gratitude to Chris Yao. We also thank the editorial staff of the PUMP Journal of Undergraduate Research for numerous suggestions that greatly improved this manuscript. This project was funded by NSF Grant DMS2341670.
@@ 1. Introduction
Example 1.1.
Suppose we have 4 dice, each with 6 faces. We roll all of them on the first turn: $3224$. We remove from the game the fourth die. On the second turn, roll all three remaining dice: $363$. Now we remove from the game the first and third dice. On the third turn, roll the remaining die: $4$. We roll that die again on the fourth turn: $1$. The game ends in four turns.
@@
In the second section, we write some initial definitions and derive the finite number of possible endings. In the third and fourth sections, we find two different formulations for the expected value and variance of the number of turns needed to finish the game. In particular, in the third section, we imagine that each die is rolled individually until a successful roll is achieved (which would remove it from the game). Thus, the problem is equivalent to studying the maximum of a sequence of i.i.d. geometric random variables, and we derive the non-recursive formulas in Theorems 3.2 and 3.3. In the fourth section, we imagine that, at each turn, all of the dice are rolled simultaneously, leading to a Markov chain interpretation and the recursive formulas in Theorems 4.1 and 4.2. In the fifth section, we prove some bounds for both the expected value and variance of the game.
@@ 2. Initial Considerations
Given $k$ dice, each with $s$ faces, a turn is formally defined as a finite sequence of $k$ numbers, each between $1$ and $s$.
@@
Let us consider Example 1.1. The signature of that game is $4331$.
@@ 3.1. Formalization
Now we imagine that each die is rolled singly until achieving a successful roll (which would remove it from the game). The probability that a die with $s$ faces, considered individually, is eliminated at each turn is constant and equal to $p=1/s$. This probability is independent of the outcomes of the other dice or the number of dice being rolled (since the faces are equiprobable). Thus, if we define $Z_{i}^{s}$ for $i=1,2,\dots,n$ the random variable for the turn in which die $i$ with $s$ faces is removed from the game, we observe that each $Z_{i}^{s}$} follows a geometric distribution with parameter $p$. Therefore, if, as before, we define $Y_{n}^{s}$ as the random variable for the number of turns to finish the game rolling $n$ dice with $s$ faces, we note that $Y_{n}^{s}$ corresponds to the turn at which the last die is eliminated, and thus
@@ Appendix B Equivalency
Note that we subtracted $1$ to correct for the term $n=0$, since $x_{0}$ does not satisfy Equation B.1. Using Lemma B.1, with $a_{j}=q^{j}x_{j}$ and $b_{n-j}=p^{n-j}$, we have

### v3

@@ Abstract.
We thank all of Polymath Jr’s staff, whose work has been key to the writing of this paper. In particular, we extend our gratitude to Chris Yao. We also thank the referee who reviewed our submission to the PUMP Journal of Undergraduate Research for numerous suggestions that greatly improved this manuscript. Lastly, we thank the editorial staff at the PUMP Journal of Undergraduate Research for their work on the journal. This project was funded by NSF Grant DMS2341670.
@@ 1. Introduction
Example $1.$
Suppose we have 4 dice, each with 6 faces. We roll all of them on the first turn: $3224$. We remove from the game the fourth die. On the second turn, roll all three remaining dice: $363$. Now we remove from the game the first and third dice. On the third turn, roll the remaining die: 44. We roll that die again on the fourth turn: 11. The game ends in four turns.
@@
In the second section, we write some initial definitions and derive the finite number of possible endings. In the third and fourth sections, we find two different formulations for the expected value and variance of the number of turns needed to finish the game. In particular, in the third section, we imagine that each die is rolled individually until a successful roll is achieved (which would remove it from the game). Thus, the problem is equivalent to studying the maximum of a sequence of i.i.d. geometric random variables, and we derive the non-recursive formulas in Theorems 3.2 and $3.$ In the fourth section, we imagine that, at each turn, all of the dice are rolled simultaneously, leading to a Markov chain interpretation and the recursive formulas in Theorems 4.1 and 4.2. In the fifth section, we prove some bounds for both the expected value and variance of the game.
@@ 2. Initial Considerations
Given $k$ dice, each with $s$ faces, a turn is formally defined as a finite sequence of $k$ numbers, each between 11 and $s$.
@@
Let us consider Example $1.$ The signature of that game is $4331$.
@@ 3.1. Formalization
Now we imagine that each die is rolled singly until achieving a successful roll (which would remove it from the game). The probability that a die with $s$ faces, considered individually, is eliminated at each turn is constant and equal to $p=1/s$. This probability is independent of the outcomes of the other dice or the number of dice being rolled (since the faces are equiprobable). Thus, if we define $Z_{i}^{s}$ for $i=1,2,\dots,n$ the random variable for the turn in which die $i$ with $s$ faces is removed from the game, we observe that each $Z_{i}^{s}$} follows a geometric distribution with parameter pp. Therefore, if, as before, we define $Y_{n}^{s}$ as the random variable for the number of turns to finish the game rolling $n$ dice with $s$ faces, we note that $Y_{n}^{s}$ corresponds to the turn at which the last die is eliminated, and thus
@@ Appendix B Equivalency
Note that we subtracted 11 to correct for the term $n=0$, since $x_{0}$ does not satisfy Equation B.1. Using Lemma B.1, with $a_{j}=q^{j}x_{j}$ and $b_{n-j}=p^{n-j}$, we have